# Learning Top-$k$ Subtask Planning Tree based on Discriminative Representation Pre-training for Decision Making

### v1

@@
Abstract.
Many complicated real-world tasks can be broken down into smaller, more manageable parts, and planning with prior knowledge extracted from these simplified pieces is crucial for humans to make accurate decisions. However, replicating this process remains a challenge for AI agents and naturally raises two questions: How to extract discriminative knowledge representation from priors? How to develop a rational plan to decompose complex problems? Most existing representation learning methods employing a single encoder structure are fragile and sensitive to complex and diverse dynamics. To address this issue, we introduce a multiple-encoder and individual-predictor regime to learn task-essential representations from sufficient data for simple subtasks. Multiple encoders are able to extract adequate task-relevant dynamics without confusion, and the shared predictor can discriminate task’s characteristics. We also use the attention mechanism to generate a top-k subtask planning tree, which customizes subtask execution plans in guiding complex decisions on unseen tasks. This process enables forward-looking and globality by flexibly adjusting the depth and width of the planning tree. Empirical results on a challenging platform composed of some basic simple tasks and combinatorially rich synthetic tasks consistently outperform some competitive baselines and demonstrate the benefits of our design.
Key words and phrases:
@@ Abstract.
Task representation, Task Planning, Reinforcement learning
1. Introduction
Figure 1. Taking two scenarios in Overcooked AI Carroll et al. (2019) for example, the top denotes cutting food, and the bottom is serving customers. Cutting food should take command of solving a sequence of subtasks: prepare the chopping board, pick up an onion, and use a knife to cut. The execution of serving customers should perform: wash dishes, dish up, and deliver.
@@ 1. Introduction
Decomposing complex problems into smaller, and more manageable subtasks with some priors and then solving them step by step is one of the critical human abilities. In contrast, reinforcement learning (RL) Schulman et al. (2017); Haarnoja et al. (2018); Konda and Tsitsiklis (1999); Babaeizadeh et al. (2016) agents typically learn from scratch and need to collect a large amount of experience, which is expensive, especially in the real world. An efficient solution is to utilize subtask priors to extract sufficient knowledge representations and transduce them into decision conditions to assist policy learning. Many real-world challenging problems with different task structures can be solved with the subtask-conditioned reinforcement learning (ScRL) regime, such as robot skill learning Kase et al. (2020); Mülling et al. (2013); Cambon et al. (2009); Gao et al. (2019); Wurman et al. (2008), navigation Zhu et al. (2017); Zhu and Zhang (2021); Quan et al. (2020); Gupta et al. (2022), and cooking Carroll et al. (2019); Knott et al. (2021); Sarkar et al. (2022). ScRL endows an agent with the learned knowledge to perform corresponding subtasks, and the subtask execution path should be planned in a rational manner. Ideally, the agent should master some basic competencies to finish some subtasks, and composite the existing subtasks to find an optimal decision path to solve previously unseen and complicated problems. For example, if the final goal is to perform an unseen task like cutting onions, the agent should be able to prepare a chopping board, pick up an onion, and use a knife. For a clear understanding, we visualize two different task structures by recombining subtask sequences, as shown in figure 1.
There are two challenges in ScRL. A fundamental question is how to construct proper knowledge representations about subtasks to transfer to more complicated tasks. The content and structure of knowledge representation are vital to the reasoning of the decision-making, since they refer to functional descriptions of different subtasks to reach the final goal. Another is how to utilize the learned knowledge to assist policy training. Most previous works Netanyahu et al. (2022); Yuan and Lu (2022); Jang et al. (2022); Ryu et al. (2022) mainly focus on zero-shot generalization on unseen tasks of the same scale, instead of recombining prior knowledge to plan more complicated tasks for guiding policy learning.
To address the above issues, our key idea is to construct proper knowledge representations, recombine existing knowledge, and apply it in policy training to improve training efficiency and model performance. We first introduce a learning regime based on multiple-encoder and individual-predictor to extract distinguishable subtask representations with task-essential properties. This module aims to exploit underlying task-specific features and construct proper knowledge representations using contrastive learning and model-based dynamic prediction. Each encoder corresponds to each subtask, which alleviates the disturbing burden of other subtasks on representation extraction to some extent. Next, we propose a construction algorithm for generating a top-$k$ subtask planning tree to customize the subtask execution plan for the agent. Then we use the subtask-conditioned reinforcement learning framework to obtain the decision policy. The top-$k$ subtask planning tree can grow to any width and depth for solving complicated tasks of different scales. Long-term planning for the future can prevent the agent from moving towards local optima as far as possible.
The contributions of this paper can be summarized as follows.
@@
  • We introduce a multiple-encoder and individual-predictor learning regime to learn distinguishable subtask representations with task-essential properties.
@@
  • Empirical evaluations on several challenging navigation tasks show that our method can achieve superior performance and obtain meaningful results consistent with intuitive expectations.
2. Related Work
Task representation
One line of research, including HiP-BMDP
@@ 2. Related Work
Zhang et al. (2020), MATE Schäfer et al. (2022), MFQI D’Eramo et al. (2020), CARE Sodhani et al. (2021), CARL Benjamins et al. (2021), etc, is designed to achieve fast adaptation by aggregating experience into a latent representation on which policy is conditioned to solve the zero/few-shot generalization in multi-task RL. Another line, such as CORRO Yuan and Lu (2022), MerPO Lin et al. (2021), FOCAL Li et al. (2020), FOCAL++ Li et al. (2021), etc, mainly focuses on learning the task/context embeddings with fully offline meta-RL to mitigate the distribution mismatch of behavior policies in training and testing. Among all the aforementioned approaches, this paper is the most related to the latter. Part of our work is devoted to finding the proper subtask representations using offline datasets to assist online policy learning.
Task planning
@@
Task planning is essentially breaking down the final goals into a sequence of subtasks. Based on the style of planning, we classify the existing approaches into three categories: planning with priors, planning on latent space, and planning with implicit knowledge. First, planning with priors will generate one or more task execution sequences according to the current situation and the subtask function prior, such as Kuffner and LaValle (2000); Liang et al. (2022); Garrett et al. (2021); Okudo and Yamada (2021), but is limited by subtasks divided by human labor. Second, planning on latent space Jurgenson et al. (2020); Pertsch et al. (2020); Zhang et al. (2021); Ao et al. (2021) endows agents with the ability to reason over an imagined long-horizon problem, but it is fundamentally bottlenecked by the accuracy of the learned dynamics model and the horizon of a task. Third, some works like Kaelbling and Lozano-Pérez (2017); Yang et al. (2022, 2020), utilize the modular combination for implicit knowledge injection to reach indirect task guidance and planning, which lack explainability. Our work aims to obtain proper knowledge representation extracted from the subtask itself, and generate a rational and explainable subtask tree including one or more task execution sequences.
Subtask-conditioned RL (ScRL)
ScRL requires the agent to make decisions according to different subtasks, which has been studied in a number of prior works Sohn et al. (2018); Sutton et al. (2022); Mehta et al. (2005); Nasiriany et al. (2019); Chane-Sane et al. (2021); Ghosh et al. (2018); Zhang et al. (2022). However, the existing works mainly solve the short-horizon subtask and rarely focus on the recombination of the learned subtask to solve the complicated long-horizon missions. We propose an explicit subtask representation and scheduling framework to provide long-horizon planning and guidance.
3. Methodology
In this section, we first describe our problem setup. Then we introduce how to learn distinguishable subtask representations as the cornerstone. Next, we propose to construct a top-$k$ subtask planning tree to customize subtask execution plans for the agent. After these, we introduce the policy learning assisted by the tree for decision making.
3.1. Problem Setup
Temporal Subtask-conditioned Markov Decision Process (TSc-MDP)
@@ 3.1. Problem Setup
We consider augmenting a finite-horizon Markov decision process Howard (1960) $(\mathcal{S},\mathcal{A},p,r,{\rho_{0}},\gamma)$ with an extra tuple ($(\{{\mathcal{T}^{i}}\}_{i=0}^{n},$ $sp,$ ${\rho_{s}})$ as the temporal subtask-conditioned Markov decision process, where $\mathcal{S}$, $\mathcal{A}$ denote the state space and the action space, respectively. $\{{\mathcal{T}^{i}}\}_{i=0}^{n}$ is the subtask set and $n$ is the number of subtasks. $p(s_{t+1},\mathcal{T}_{t+1}^{i}|s_{t},\mathcal{T}_{t}^{i},a_{t})$ is the dynamics function that gives the distribution of the next state $s_{t+1}$ and next subtask $\mathcal{T}_{t+1}^{i}$ at state $s_{t}$ to execute the subtask $\mathcal{T}_{t}^{i}$ and take action $a_{t}$. $r:\mathcal{S}\times\mathcal{T}\times\mathcal{A}\to\mathbb{R}$ denotes the reward function. $\rho_{0}$ and $\rho_{s}$ denote the initial state distribution and the subtask distribution. $\gamma\in(0,1)$ is the discount factor. $sp$ is the temporal subtask planner that generates subtask sequences at a coarse time scale. Formally, the optimization objective is to find the optimal policy $\pi$ and the subtask planner $sp$ to maximize the expected cumulative return:
@@
where $\tau$ is the history transitions.
3.2. Distinguishable Subtask Representations
Data Collection.
@@ 3.2. Distinguishable Subtask Representations
Due to the time-consuming and often ineffective training for complex tasks, directly collecting expert data is intractable. Therefore, we disassemble the complex task into a set of multiple simple tasks, denoted as $\mathcal{T}:\{\mathcal{T}^{1},\mathcal{T}^{2},...,\mathcal{T}^{n}\}$. Simple tasks can easily be trained from scratch to obtain optimal policies. For collecting diverse and relevant data and eliminate the effect of the behaviour policies, we collect 1000 trajectories separately from the optimal, the intermediate trained, and the random policy for each subtask. The collected data format for subtask $\mathcal{T}^{i}$ can be denoted as $\{(s_{t},a_{t},r_{t},s_{t+1})_{t=1}^{L}\}^{i}$, where $L$ is the trajectory length. The data collection statistics are shown in Appendix A.1.
Representation Pre-training.
@@
In reinforcement learning domain, the environmental system is characterized by a transition function and a reward function. Therefore, what we need to do is to use the collected trajectories to mine the essential properties of subtasks and make proper representations from the perspective of the transition and the reward function. To learn accurate and discriminative subtask representations, we introduce a multiple-encoder and individual-predictor learning regime to extract task-essential properties, as shown in Figure 2.
@@
Figure 2. The multiple-encoder and individual-predictor learning regime. Each subtask encoder $\mathcal{E}^{i}$ is used to encode the state-action pair $(s_{t},a_{t})^{i}\in\mathcal{T}^{i}$ and output the compact subtask representation $z_{t}^{i}$. The shared subtask predictor is used to predict the reward $r_{t}$ and the next state $s_{t+1}$. The contrastive and prediction losses are designed to learn the many to one framework.
@@
There are two design principles: i) Each subtask has its own different responsibilities and skills, so multiple encoders can boost the diversity among them. On the other hand, multiple subtask encoders, which each corresponds to one subtask, can effectively avoid mapping two state-action pairs from different subtasks with various reward and transition distributions to similar abstractions. ii) Figuring out the essential properties of transition and reward function is the key to obtaining accurate and generalizable subtask representations, so we design an auxiliary prediction task to simulate the reward and transition distributions. The shared subtask predictor is used to predict the reward and the next state so as to implicitly learn the world model for a better understanding of the accurate essential properties of the subtasks.
@@
After obtaining the latent variables of different subtasks, for better distinguishment, a contrastive loss is applied, where samples from the other subtasks act as negative pairs. First, we define $x=(s,a)$, $x^{+}=(s^{+},a^{+})$, and $x^{-}=(s^{-},a^{-})$ as the current, the positive, and the negative state-action pairs, respectively. Given the current subtask id $i$, we have the latent variables, denoted as $q={\mathcal{E}^{i}}(s,a)$, $k^{+}={\mathcal{E}^{i}}(s^{+},a^{+})$, $k^{-}={\mathcal{E}^{j}}(s^{-},a^{-})$, and $j$ is the other subtask ids except $i$. Therefore, the contrastive learning objective Oord et al. (2018) for the multiple subtask encoders is:
@@
As for the shared subtask predictor $f_{pre}$, we use the mean squared error (MSE) loss to calculate the mean overseen data of the squared differences between true and prediction as follows.
${\mathcal{L}_{p}}={\lambda_{r}}||\hat{r}-r||_{2}^{2}+{\lambda_{s}}||\hat{s}^{\prime}-s^{\prime}||_{2}^{2}\ ,$ (3)
where $\hat{r},\hat{s}^{\prime}=f_{pre}(z)$ is the prediction of the immediate reward $r$ and the next state $s^{\prime}$. $\lambda_{r}$ and $\lambda_{s}$ are positive scalars that tune the relative weights between the loss terms.
With contrastive learning for pushing inter-tasks and pulling intra-task and dynamics prediction for mining nature properties, the total loss is summarized as follows.
@@
After finishing the first step toward compact representations for subtasks, the subtask tree can be generated.
3.3. Top-$k$ Subtask Planning Tree Generation
Figure 3. The left part is a top-$2$ subtask tree [^2] Note: The probability sum is not equal to 1 because we sample 2 subtasks from the original distribution and ignore other subtasks.which is generated by the process of the right part.
In real human society, effective decision-making always relies on good judgment of the current state, and future prediction and imagination. Enlightened by this ability of humanity, we should generate a top-$k$ subtask tree to assist the policy training. Due to the long horizon and sparse feedback of complicated tasks, we adjust the subtask tree generation during rollout instead of generating a specified tree for one task. As shown in figure 2, we generate a relatively short-horizon subtask tree with some inference of dynamics in a coarse-to-fine manner and a new subtask tree will regenerate once the current tree has been executed to a leaf node. Next, we describe the generation process in detail as follows.
First, we provide the detailed construction procedure of the subtask tree in a top-down manner. Here, we have a query encoder $\mathcal{E}^{q}$ and the $n$ pre-trained encoder $\{\mathcal{E}^{i}\}^{n}_{i=1}$ corresponding to each subtask defined in section 3.2.
Node-edge construction process.
At the current timestep $t$, due to the limitation that we can only touch the last action $a_{t-1}$, we compute the query vector $v^{q}$ as the anchor to retrieve the most similar subtask.
${v^{q}}={{{{\mathcal{E}}}}^{q}}({s_{t-1}},{a_{t-1}}).$ (5)
Next, the subtask embeddings from the pre-trained encoder can be obtained as follows.
$\{{v^{i}}\}_{i=1}^{n}=\{{{\mathcal{E}}^{i}}({s_{t-1}},{a_{t-1}})\}_{i=1}^{n}.$ (6)
We utilize the attention mechanism Bahdanau et al. (2015) to compute the correlation score between the query vector $v^{q}$ and the pre-trained subtask embeddings $\{{v^{i}}\}_{i=1}^{n}$. Then the normalized coefficient $\alpha_{q,i}$ is defined:

### v2

@@
Abstract
Decomposing complex real-world tasks into simpler subtasks and devising a subtask execution plan is critical for humans to achieve effective decision-making. However, replicating this process remains challenging for AI agents and naturally raises two questions: (1) How to extract discriminative knowledge representation from priors? (2) How to develop a rational plan to decompose complex problems? To address these issues, we introduce a groundbreaking framework that incorporates two main contributions. First, our multiple-encoder and individual-predictor regime goes beyond traditional architectures to extract nuanced task-specific dynamics from datasets, enriching the feature space for subtasks. Second, we innovate in planning by introducing a top-$K$ subtask planning tree generated through an attention mechanism, which allows for dynamic adaptability and forward-looking decision-making. Our framework is empirically validated against challenging benchmarks BabyAI including multiple combinatorially rich synthetic tasks (e.g., GoToSeq, SynthSeq, BossLevel), where it not only outperforms competitive baselines but also demonstrates superior adaptability and effectiveness in complex task decomposition.
keywords:
@@ Abstract
Reinforcement Learning, Representation Learning, Subtask Planning, Task Decomposition, Pretraining
[…]
1 Introduction
@@
Decomposing complex problems into smaller, and more manageable subtasks with some priors and then solving them step by step is one of the critical human abilities. In contrast, reinforcement learning (RL) schulman2017ppo ; haarnoja2018sac ; konda1999actor_critic ; dang2022dynamic agents typically learn from scratch and need to collect a large amount of experience, which can be costly, particularly in real-world scenarios. To emulate the efficient decision-making process of humans, one solution is to leverage subtask priors to extract useful knowledge representations and transduce them into decision conditions to assist policy learning. Many real-world challenging problems with different task structures can be solved with the subtask-conditioned reinforcement learning (ScRL) regime, such as robot skill learning kase2020transferable ; mulling2013learning ; cambon2009hybrid ; li2021skill ; wurman2008coordinating , navigation zhu2017target ; zhu2021deep ; quan2020novel ; gupta2022intention , and cooking carroll2019utility ; knott2021evaluating ; sarkar2022pantheonrl . ScRL endows an agent with the learned knowledge to perform corresponding subtasks, and the subtask execution path should be planned in a rational manner. Ideally, the agent should master some basic competencies to finish some subtasks and composite the existing subtasks to find an optimal decision path to solve previously unseen and complicated problems. Taking the task of cutting onions for example, consists of multiple interdependent subtasks, including preparing a chopping board, procuring an onion, and utilizing a knife. For clarity, we visualize two different task structures by recombining subtask sequences, as shown in Fig. 1. It is crucial to integrate the subtasks in a sequential manner to accomplish the ultimate goal of onion cutting.
Figure 1: Taking two scenarios in Overcooked AI carroll2019utility as examples, the top denotes cutting food, and the bottom is serving customers. Cutting food should take command of solving a sequence of subtasks: prepare the chopping board, pick up an onion, and use a knife to cut. The execution of serving customers should be performed: wash dishes, dish up, and deliver.
However, ScRL faces two major challenges. The first challenge involves the construction of proper knowledge representations for subtasks, which can facilitate the transfer to more complex tasks. The content and structure of these representations are critical to decision-making, as they offer functional descriptions of different subtasks leading to the completion of the final goal. Addressing this challenge requires a nuanced understanding of the dynamics of each subtask, which we propose can be learned from their trajectories. To operationalize this, we introduce a novel learning regime based on multiple encoders and an individual predictor aimed at extracting distinguishable subtask representations that encapsulate these dynamics along with other task-essential properties. This framework utilizes contrastive learning and model-based dynamic prediction to capture these features effectively. Importantly, each encoder is designated to a specific subtask, minimizing interference from other subtasks and enhancing the quality of the extracted representations.
Another key challenge in ScRL is how to leverage learned knowledge to improve policy training. While prior works netanyahu2022discovering ; kwan2023survey ; jang2022bc ; ryu2022remax have primarily focused on zero-shot generalization to unseen tasks of the similar complexity, little attention has been paid to the recombination of prior knowledge to plan more intricate tasks that can guide policy learning. This narrow focus limits the agent’s ability to adapt to more complex or unfamiliar scenarios. To address this limitation, we introduce a novel construction algorithm that generates a top-$K$ subtask planning tree, customizing the execution plan for each subtask. This approach not only tailors the learning process to the specific needs of the agent but also allows for the incorporation of learned knowledge into the decision-making policy. Unlike traditional methods, the top-$K$ subtask planning tree can be expanded to any width and depth, providing a flexible and scalable solution for complex tasks. Our top-$K$ subtask planning tree facilitates a more comprehensive view by enabling the agent to look ahead and consider a range of possible future subtasks and their implications. By doing so, the agent can more accurately weigh the immediate rewards against potential future gains or losses. This long-term outlook minimizes the chances of the agent getting stuck in locally optimal solutions that are suboptimal in the broader context. In essence, long-term planning adds a layer of foresight that contributes to optimizing the overall decision-making process, making the agent better equipped to handle complex tasks.
In summary, the contributions of this paper can be summarized in threefold.
@@
  • We introduce a multiple-encoder and individual-predictor learning regime to learn distinguishable subtask representations with task-essential properties, which enables the agent to acquire generalized subtask knowledge.
@@
  • Empirical evaluations on several challenging navigation tasks (e.g., GoToSeq, SynthSeq, BossLevel) show that our method can achieve superior performance and obtain meaningful results consistent with intuitive expectations.
2 Related Work
@@
Task representation. One line of research, including HiP-BMDP zhang2020learning , MATE schafer2022learning , MFQI d2020mfqi , CARE sodhani2021care , CARL benjamins2021carl , etc, are designed to achieve fast adaptation by aggregating experience into a latent representation on which policy is conditioned to solve the zero/few-shot generalization in multi-task RL. Another line, such as CORRO yuan2022corro , MerPO lin2021MerPO , FOCAL li2020focal , CMT lin2022contextual , HS-OMRL zhao2023context , etc, mainly focus on learning the task/context embeddings with fully offline meta-RL to mitigate the distribution mismatch of behavior policies in training and testing. Among all the aforementioned approaches, this paper is the most related to the latter. Part of our work is devoted to finding the proper subtask representations using offline datasets to assist online policy learning.
@@
Task planning. Task planning essentially breaks down the final goals into a sequence of subtasks. Based on the style of planning, we classify the existing approaches into four categories: planning with priors, planning on latent space, and planning with implicit knowledge. First, planning with priors will generate one or more task execution sequences according to the current situation and the subtask function prior, such as kuffner2000rrt ; liang2022search ; garrett2021integrated ; okudo2021online , but is limited by subtasks divided by human labor. Second, planning on latent space jurgenson2020sub ; pertsch2020long ; zhang2021world ; ao2021co ; zhang2018learning endows agents with the ability to reason over an imagined long-horizon problem, but it is fundamentally bottlenecked by the accuracy of the learned dynamics model and the horizon of a task. Third, some works such as kaelbling2017learning ; yang2022ldsa ; yang2020multi ; ruan2023learning , utilize the modular combination for implicit knowledge injection to achieve indirect task guidance and planning, which lack explainability. Finally, recent developments in the field have seen the emergence of work such as TPTU ruan2023tptu , ProgPrompt singh2023progprompt , Plan4mc yuan2023plan4mc , and Voyager wang2023voyager , which leverage large language models (LLMs) for task planning. However, a significant drawback of using LLMs is their substantial computational cost. Our work aims to obtain proper knowledge representation extracted from the subtask and generate a rational subtask tree including one or more task execution sequences.
Subtask-conditioned RL (ScRL). ScRL requires the agent to make decisions according to different subtasks. A line of work sohn2018hierarchical ; sutton2022reward ; mehta2005multi ; nasiriany2019planning ; chane2021goal ; ghosh2018learning ; zhang2022multi focuses on leveraging the prefined or learned abstract subtasks, which lack transparency and have difficulty solving long-horizon problems. There is some work that tends to develop some explainable methods. From the perspective of the neuro symbol, BPNS silverlearning learns neuro-symbolic skills and uses heuristic search A* to guide decision-making. Pulkit et al. verma2022discovering propose learning user-interpretable capacities from manually collected trajectories and the user’s vocabulary. However, how to exploit these subtasks to make long-term planning is not considered, which is very unfavorable for complex tasks. However, the existing works mainly solve the short-horizon subtask and rarely focus on the recombination of the learned subtask to solve the complicated long-horizon missions. We propose an explicit subtask representation and scheduling framework to provide long-horizon planning and guidance.
3 Problem Statement
In this section, we will commence by introducing some fundamental concepts, followed by a formal definition of the temporal subtask-conditioned Markov decision process.
3.1 Definitions
In this paper, we investigate how to solve a complicated task that contains some simpler parts by representation and planning. Some fundamental concepts are explained as follows.
  • Subtask. A subtask is a smaller, more manageable component of a complex task. The completion of the task often requires an orderly combination of multiple subtasks.
  • Subtask representation. Subtask representation refers to the encoding of the essential information that characterizes a subtask, which is typically learned from the data and used by an agent to facilitate the identification and execution of subtasks during a complex task.
  • Subtask planning. In reinforcement learning, subtask planning is often formalized as a sequential decision-making problem, where the agent chooses a sequence of subtasks to perform, in order to maximize the cumulative reward over the entire task.
3.2 Problem Setup
@@
Temporal Subtask-conditioned Markov Decision Process (TSc-MDP). We consider augmenting a finite-horizon Markov decision process howard1960dynamic $(\mathcal{S},\mathcal{A},p,r,{\rho_{0}},\gamma)$ with an extra tuple ($(\{{\mathcal{T}^{i}}\}_{i=1}^{n},$ $sp,$ ${\rho_{s}})$ as the temporal subtask-conditioned Markov decision process, where $\mathcal{S}$, and $\mathcal{A}$ denote the state space and the action space, respectively. $\{{\mathcal{T}^{i}}\}_{i=1}^{n}$ is the subtask set and $n$ is the number of subtasks. $p(s_{t+1},\mathcal{T}_{t+1}^{i}|s_{t},\mathcal{T}_{t}^{i},a_{t})$ is the dynamics function that gives the distribution of the next state $s_{t+1}$ and next subtask $\mathcal{T}_{t+1}^{i}$ at state $s_{t}$ to execute the subtask $\mathcal{T}_{t}^{i}$ and take action $a_{t}$. $r:\mathcal{S}\times\mathcal{T}\times\mathcal{A}\to\mathbb{R}$ denotes the reward function. $\rho_{0}$ and $\rho_{s}$ denote the initial state distribution and the subtask distribution, respectively. $\gamma\in(0,1)$ is the discount factor. $sp$ is the temporal subtask planner that generates subtask sequences at a coarse time scale. Formally, the optimization objective is to find the optimal subtask planner $sp$ and policy $\pi$ to generate the orderly subtask plans and then make a decision for maximizing the expected cumulative return:
@@
where $\tau$ is the history transition. In our setting, we focus on how to extract the subtask representation and then devise a rational subtask execution plan for the AI agent to make efficient decisions.
4 Methodology
In this section, we begin by articulating the underlying motivation that guided the development of our approach. Following this, we elaborate on our unique method for learning distinguishable subtask representations, serving as the foundational element of our algorithm. Subsequently, we introduce our innovative construction of a top-$K$ subtask planning tree, specifically designed to customize the agent’s subtask execution plans. Finally, we discuss how this planning tree facilitates and enhances policy learning for effective decision-making.
4.1 The Motivation for the Proposed Framework
Current reinforcement learning methods often face two major challenges when tackling complex tasks: (a) Suboptimal Knowledge Representation: Traditional single-encoder models struggle to capture the unique dynamics of individual subtasks. This leads to poor generalization and inefficiency in learning complex tasks. (b) Lack of Long-term Planning: Current RL algorithms are often myopic, focusing only on immediate rewards. This short-sightedness often leads to suboptimal policies, as agents can get stuck in local optima without considering long-term consequences.
To address the first challenge of suboptimal knowledge representation, we should focus on:
  • Specialization for Each Subtask: We propose a multiple-encoder architecture to provide specialized learning and representation for each subtask, thereby tackling the inefficiency caused by one-size-fits-all models.
  • Task-Essential Feature Extraction: An individual predictor accompanying each encoder is designed to capture the unique, task-relevant features, enabling better discrimination between different subtasks.
To address the second challenge of the lack of long-term planning, some factors should be taken into consideration:
  • Enabling Forward-Looking Decisions: We introduce the concept of a top-$K$ subtask planning tree to guide the agent’s policy. This allows for longer-term planning by offering a roadmap of potential future subtasks.
  • Dynamic Scalability: The top-$K$ subtask planning tree is not fixed in its breadth or depth, offering flexibility in addressing tasks of varying complexity.
  • Global Optima over Local Optima: By giving the agent the ability to consider a range of future possibilities, the planning tree aids the agent in making decisions that are globally optimal rather than just locally optimal.
By addressing these identified challenges, our proposed framework aims to fill the existing gaps in the domain of RL, providing a more effective and comprehensive solution for complex decision-making tasks. In the subsequent sections, we will give detailed descriptions about three core components of our methodology: (a) the extraction of distinguishable subtask representations through a multiple-encoder and individual-predictor regime, (b) the generation of a top-$K$ subtask planning tree, and (c) the formulation of a tree-based policy for better decision-making.
4.2 Distinguishable Subtask Representations
@@
Data Collection. Due to the time-consuming and often ineffective training for complex tasks, directly collecting expert data is intractable. Therefore, we disassemble the complex task into a set of multiple simple tasks, denoted as $\mathcal{T}:\{\mathcal{T}^{1},\mathcal{T}^{2},...,\mathcal{T}^{n}\}$. Simple tasks can easily be trained from scratch to obtain optimal policies. To collect diverse and relevant data and eliminate the effect of the behavior policies, we separately collect 3000 trajectories for each subtask, of which 1/3 is collected from the converged PPO schulman2017ppo , 1/3 is collected from the intermediate trained checkpoint of PPO, and the last 1/3 is from the random policy. The data statistics are shown in Table 1. The collected data format for subtask $\mathcal{T}^{i}$ can be denoted as $\{(s_{t},a_{t},r_{t},s_{t+1})_{t=1}^{L}\}^{i}$, where $L$ is the trajectory length.
Subtask #Total Episodes #Total Transitions
PickupLoc 3000 210375
PutNext 3000 559201
Goto 3000 218191
Open 3000 149349
Table 1: The data collection statistics
@@
Representation Pretraining. In reinforcement learning, the environment is modelled by a transition function and a reward function. We use the collected trajectories to extract the essential properties of subtasks and create proper representations based on these functions. To achieve this, we introduce a multiple-encoder and individual-predictor learning regime that can effectively capture the task-essential properties, as shown in Fig. 2.
@@
Figure 2: The multiple-encoder and individual-predictor learning regime. Each subtask encoder $\mathcal{E}^{i}$ encodes the state-action pair $(s_{t},a_{t})^{i}\in\mathcal{T}^{i}$ and outputs the compact subtask representation $z_{t}^{i}$ at each time step. The shared subtask predictor predicts the reward $r_{t}$ and the next state $s_{t+1}$. The contrastive and prediction losses are designed to train the subtask encoders and the predictor end-to-end.
@@
Our framework incorporates two design principles. First, we use multiple encoders to enhance diversity among subtasks, as each subtask has its own unique responsibilities and skills. This approach also ensures that state-action pairs from different subtasks are not mapped to similar abstractions. Second, we introduce an auxiliary prediction task to simulate reward and transition distributions, allowing us to learn accurate and generalizable subtask representations. By predicting the reward and next state, our shared subtask predictor implicitly learns the world model, facilitating a better understanding of the essential properties of the subtasks.
@@
For better distinguishment, the contrastive loss is applied. Similar to  yuan2022corro ; oord2018representation , we give a definition of anchors, positive, and negative samples. Using the latent variable $q={\mathcal{E}^{i}}(s,a)$ from the subtask $\mathcal{T}^{i}$ as an anchor, the $k^{+}={\mathcal{E}^{i}}(s^{+},a^{+})$ and $k^{-}={\mathcal{E}^{j}}(s^{-},a^{-})$ respectively constitute positive and negative instances, where $(s^{+},a^{+})$ and $(s^{-},a^{-})$ are sampled from the subtask $\mathcal{T}^{i}$ and other subtasks $\mathcal{T}^{j},j\neq i$. Thus, the contrastive objective oord2018representation is:
@@
Then adopting mini-batch sampling for training, we can optimize the contrastive objective to extract the essential variance of transition functions across different tasks and mine the shared and compact features of the same subtasks.
As for the shared subtask predictor $f_{pre}$, we use the mean squared error (MSE) loss to calculate the mean squared differences between true and prediction as follows.
${\mathcal{L}_{p}}={\lambda_{r}}\|\hat{r}-r\|_{2}^{2}+{\lambda_{s}}\|\hat{s}^{\prime}-s^{\prime}\|_{2}^{2}\ ,$ (3)
where $\hat{r},\hat{s}^{\prime}=f_{pre}(z)$ is the prediction of the immediate reward $r$ and the next state $s^{\prime}$. $\lambda_{r}$ and $\lambda_{s}$ are positive scalars that tune the relative weights.
Thus, the proposed framework includes a total loss that combines the contrastive loss, which encourages similar subtask representations for state-action pairs from the same subtask and distinct representations for those from different subtasks, and the dynamics prediction loss, which aims to capture the essential properties of the transition and reward functions, which is summarized as follows.
@@
In summary, by disassembling complex tasks into simpler, manageable pieces, our approach enables the collection of diverse, various-quality data for each subtask. These data are then channeled through a novel multi-encoder and single predictor paradigm, designed to adeptly grasp the unique dynamics of each subtask, laying a strong foundation for advanced planning and decision-making modules. The representation pretraining fortifies each encoder’s specialization in their respective subtasks, ensuring enhanced, distinct learning and representation. This comprehensive approach guarantees a robust and compact subtask representation for downstream decision-making. After finishing the first step toward compact representations for subtasks, the subtask tree can be generated for the planning and decision-making process.
4.3 Top-K Subtask Planning Tree Generation
Effective decision-making in human society often involves considering the current state, predicting the future, and imagining various outcomes. Inspired by this, we propose a method to generate a top-$K$ subtask tree to assist with policy training. To handle long horizons and sparse feedback in complex tasks, we dynamically adjust the subtask tree generation during one episode instead of using a predefined tree. Specifically, we generate a subtask tree in a coarse-to-fine manner with some inference of dynamics, as illustrated in Fig. 3a. Then we regenerate a new subtask tree once the current tree has been executed to a leaf node. Next, we describe the generation process of the node and edge in detail as follows.
First, we provide the detailed node-edge construction procedure of the subtask tree in a top-down manner. Here, we have a query encoder $\mathcal{E}^{q}$ and the $n$ pretrained encoder $\{\mathcal{E}^{i}\}^{n}_{i=1}$ corresponding to each subtask defined in section 4.2.
(a) The framework
(b) The pseudo code
Figure 3: The generation of the top-$K$ subtask planning tree [^2] Note: Attention function is defined by Equation (5-8) and Planning function is defined by the $m$-step predictor $p_{m}$..
Node-edge construction. Here, we have $n$ pretrained encoder $\{\mathcal{E}^{i}\}^{n}_{i=1}$ corresponding to each subtask defined in section 4.2. We give a query encoder $\mathcal{E}^{q}$ to retrieve the subtask, which serves as the node in the tree. Given the last state-action pair $(s_{t-1},a_{t-1})$, we can compute the query vector $v^{q}$ and the subtask embeddings $\{{v^{i}}\}_{i=1}^{n}$ as follows.
${v^{q}}={{{{\mathcal{E}}}}^{q}}({s_{t-1}},{a_{t-1}}),\ \ \ \ \ \{{v^{i}}\}_{i=1}^{n}=\{{{\mathcal{E}}^{i}}({s_{t-1}},{a_{t-1}})\}_{i=1}^{n}.$ (5)
Then we utilize the attention mechanism bahdanau2015attention to compute the correlation score between $v^{q}$ and $\{{v^{i}}\}_{i=1}^{n}$. Then the normalized coefficient $\alpha_{q,i}$ is defined: